# ACC Saturator: Automatic Kernel Optimization for Directive-Based GPU Code

### v1

@@ 6.2. Bulk Load
| cg | Eigenvalue | Irregular | 16 | Ref | (> CLASS C) | 4.28s | 662.58s | 5.06s | 19.03s | 18.42s |
@@
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
7. Experimental Methodology
We implement ACC Saturator in Racket (Felleisen et al., 2015) using XcodeML (CCS), 2009) to parse and generate OpenACC/OpenMP source codes in C. We use the egg library (Willsey et al., 2021) to perform equality saturation. ACC Saturator integrates with NVHPC (NVIDIA Corporation, 2023) and GCC (GNU Project, 2023) for OpenACC/OpenMP compilation, and with Clang (The LLVM Project, 2023) for OpenMP compilation. For evaluation, we use NVHPC 22.9 with options "-O3 -gpu=fastmath -Msafeptr -(acc|mp)=gpu", GCC 12.2.0 with "-O3 -ffast-math -f(openacc|openmp)" and Clang 15.0.3 with "-O3 -ffast-math -fopenmp". Our experiments run on an NVIDIA A100-PCIE-40GB GPU with an Intel Xeon Silver 4114 CPU.
We evaluate the kernel-execution performance of ACC Saturator on two benchmark suites: NAS Parallel Benchmarks in OpenACC/C (NPB) (Xu et al., 2014) and the SPEC ACCEL benchmark suite in both OpenACC/C and OpenMP/C (SPEC) (Corporation, 2014). Tables 2 and 3 provide the detail of NPB and SPEC, respectively. To ensure adequate memory usage, we select CLASS C as the problem sizes of all NPB benchmarks (the largest size within standard test problems), SPEC uses the referential sizes (Ref) except for GCC’s OpenACC cases of ep, sp, and bt, for which we select the testing sizes (Test) due to high execution latency. NPB’s BT, CG, EP, and SP feature the same computation as SPEC’s bt, cg, ep, and csp, but the implementation of NPB is based on OpenACC’s parallel directive while that of SPEC’s OpenACC benchmarks is on the kernels directive. We report the best kernel performance of three executions. To avoid producing incorrect results, we remove the user-specific parallelism in NPB’s CG for GCC, disable the device-side reduction of GCC’s OpenACC in NPB’s LU and MG, and omit the degree specification of the worker parallelism from GCC’s NPB cases, since it surpasses GCC’s thread limit.
@@ 8. Evaluation
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
@@
Table 4 breaks down the top 10 kernels in NPB’s BT, and Figure 3 shows the speedup of each kernel in each version. The speedups of CSE+BULK and ACCSAT are similar to those of CSE and CSE+SAT, respectively. ACCSAT attains up to 2.23x and 5.08x improvements on NVHPC and GCC, respectively, resulting in up to 2.08x and 3.19x memory bandwidth. The top three kernels on NVHPC suffer performance degradation from CSE+BULK to ACCSAT, because ACCSAT spills more registers to global memory, facilitating the reuse of computation. The next three kernels execute 8.3% fewer instructions and around a 1.11x speedup compared to CSE+BULK, since our optimized code clarifies dependencies and reduces both computation and stores. We expect a total speedup of 1.22x when equality saturation is disabled when registers spill.
@@
Figure 4 depicts the speedups of SPEC benchmarks for both OpenACC and OpenMP versions. Benchmarks with names starting with "p-" indicate the OpenMP versions. On average, with OpenACC, CSE, CSE+SAT, CSE+BULK, and ACCSAT attain 1.06x, 1.04x, 1.08x, and 1.08x speedups, respectively on NVHPC, while with OpenMP, these attain 1.07x, 1.06x, 1.37x, and 1.47x speedups, respectively. OpenACC’s ostencil sees a 16.7% reduction in SM occupancy due to equality saturation, leading to decreased performance. CSE reduces memory loads by around 50% for olbm and polbm, resulting in improved performance. Bulk load significantly boosts the performance of memory/intensive benchmarks such as csp, bt, pcsp, and pbt, just as in NPB. ACCSAT attains 3.62x speedup in one part of pbt, which executes only one thread-block over nested loops, by eliminating 77.2% memory loads and 50.9% stores. Our optimization decreases operations and reorders memory accesses, being effective for both parallel and sequential iterations.
@@
9. Related Work
Maximizing architectural utilization is a significant challenge in HPC, requiring application code optimization. Compilers play a crucial role for enhancing performance by applying various code-generation techniques, enabling both generic and architecture-specific optimizations. Numerous programming models and state-of-the-art techniques, especially for GPUs, have been introduced to adapt codes to parallel architectures (Ragan-Kelley et al., 2013; NVIDIA Corporation, 2022; Stulajter et al., 2022; Paszke et al., 2019; Edwards and Sunderland, 2012). Directive-based programming models (Organization, 2011; ARB, 1997) extend sequential languages, allowing complex scientific applications to offload loop iterations to accelerators while maintaining their structures. However, such compilers often rely on the basis of sequential code generation, thus limiting optimization opportunities to general computation (Matsumura et al., 2021; Barua et al., 2018).
Several projects have explored domain-specific or architecture/specific approaches for directive-based code. The CLAW DSL (Clement et al., 2018) provides directives for grid-based algorithms, enabling target-specific optimizations such as spatial blocking while supporting OpenACC/OpenMP code generation. JACC (Matsumura et al., 2021) is an OpenACC runtime framework offering just-in-time kernel compilation with dynamic constant expansion. OptACC (Montgomery et al., 2015) performs runtime parameter searches to optimize OpenACC parallelism. CCAMP (Lambert et al., 2020) interchanges OpenACC and OpenMP, optimizing parallelization for each combination of models and architecture. Barua et al. (Barua et al., 2018) develop an automated OpenACC-kernel optimizer for maximizing ILP through loop unfolding. SAFARA (Tian et al., 2016) fully utilizes register resources to facilitate array reference reuse in OpenACC kernels.
@@ 9. Related Work
Since the introduction of the equality-saturation library egg (Willsey et al., 2021), numerous studies have leveraged it, particularly in the context of GPU computing, for accelerating deep learning applications (Yang et al., 2021; Fu et al., 2023; Shaikhha et al., 2023; Wang et al., 2020; Smith et al., 2021). These works employ rewriting rules for arithmetic expressions, abstract operations, or tensor graphs to optimize convolutions, sparse tensor algebra, or whole tensor operations. Diospyros (VanHattum et al., 2021) synthesizes efficient DSP operations from C code using equality saturation, while Gowda et al. (Gowda et al., 2022) implement a symbolic algebra system with egg for automatic parallelism assignment.
Although the initial work of equality saturation was demonstrated as a Java bytecode optimizer (Tate et al., 2009), recent works using egg focus on program synthesis and code optimization without control statements. Our work is the first to bridge the gap between user code and equality saturation by automatically extracting computation and constructing SSA information for data dependencies. This approach enables novel equality/saturation optimizations for directive-based programming, as detailed in this paper, without requiring further programmer intervention.
Several works propose innovative code optimization approaches. Ben-Nun et al. (Ben-Nun et al., 2019; Ben-Nun et al., 2023) develop data-centric flow graphs to focus on data-oriented optimizations, such as dead memory elimination. Their representation supports collective operations and macroscopic parallelization, enabling optimizations irrespective of calculation. The MLIR framework (Chelini et al., 2021) utilizes multi-level IRs for cooperative domain-specific optimizations. Ginsbach et al. (Ginsbach et al., 2018) define code patterns for performance opportunities, replacing matched code with library or DSL implementations. Additionally, machine learning is gaining popularity for automatic compiler-optimization tuning (Park et al., 2022; Trofin et al., 2021; Haj-Ali et al., 2020). Our tool employs e-graph operations to identify optimal solutions without source code analysis or abstract syntax tree transformations.
Numerous optimization techniques have been developed for GPU computing (Hijma et al., 2023). Rawat et al. (Rawat et al., 2018) reorder stencil computation using heuristic algorithms to alleviate register pressure and increase SM occupancy. Software systolic arrays (Chen et al., 2019) assign global data to GPU threads, computing results by propagating and accumulating partial results through warp-level shuffles across overlapped blocks, improving data locality without shared memory usage. Hong et al. (Hong et al., 2018) utilize lightweight kernel emulation to provide auto-tuners with performance bottleneck feedback. In contrast, ACC Saturator adopts a straightforward rewriting mechanism, enabling its application to complex programs. Moreover, our method is complementary to other code optimizations.
10. Conclusion
@@
Acknowledgement
We are funded by the EPEEC project from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 801051 and the Ministerio de Ciencia e Innovación—Agencia Estatal de Investigación (PID2019-107255GB-C21/AEI/10.13039/501100011033). We gratefully acknowledge the support of NVIDIA Solutions Lab who provided us the remote access to their GPU environment. We would like to acknowledge the NVIDIA AI Technology Center (NVAITC) Europe for their valuable help. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.

### v3

@@ VI-B Bulk Load
| cg | Eigenvalue | Irregular | 16 | Ref | (¿ CLASS C) | 4.28s | 662.58s | 5.06s | 19.03s | 18.42s |
@@
VII Experimental Methodology
We implement ACC Saturator in Racket [32] using XcodeML [33] to parse and generate OpenACC/OpenMP source codes in C. We use the egg library [21] to perform equality saturation. ACC Saturator integrates with NVHPC [34] and GCC [35] for OpenACC/OpenMP compilation, and with Clang [36] for OpenMP compilation. For evaluation, we use NVHPC 22.9 with options "-O3 -gpu=fastmath -Msafeptr -(acc|mp)=gpu", GCC 12.2.0 with "-O3 -ffast-math -f(openacc|openmp)" and Clang 15.0.3 with "-O3 -ffast-math -fopenmp". Our experiments run on an NVIDIA A100-PCIE-40GB GPU with an Intel Xeon Silver 4114 CPU.
We evaluate the kernel-execution performance of ACC Saturator on two benchmark suites: NAS Parallel Benchmarks in OpenACC/C (NPB) [30] and the SPEC ACCEL benchmark suite in both OpenACC/C and OpenMP/C (SPEC) [31]. Tables II and III provide the detail of NPB and SPEC, respectively. To ensure adequate memory usage, we select CLASS C as the problem sizes of all NPB benchmarks (the largest size within standard test problems), SPEC uses the referential sizes (Ref) except for GCC’s OpenACC cases of ep, sp, and bt, for which we select the testing sizes (Test) due to high execution latency. NPB’s BT, CG, EP, and SP feature the same computation as SPEC’s bt, cg, ep, and csp, but the implementation of NPB is based on OpenACC’s parallel directive while that of SPEC’s OpenACC benchmarks is on the kernels directive. We report the best kernel performance of three executions. To avoid producing incorrect results, we remove the user-specific parallelism in NPB’s CG for GCC, disable the device-side reduction of GCC’s OpenACC in NPB’s LU and MG, and omit the degree specification of the worker parallelism from GCC’s NPB cases, since it surpasses GCC’s thread limit.
@@ VIII Evaluation
Table IV breaks down the top 10 kernels in NPB’s BT, and Figure 3 shows the speedup of each kernel in each version. The speedups of CSE+BULK and ACCSAT are similar to those of CSE and CSE+SAT, respectively. ACCSAT attains up to 2.23x and 5.08x improvements on NVHPC and GCC, respectively, resulting in up to 2.08x and 3.19x memory bandwidth. The top three kernels on NVHPC suffer performance degradation from CSE+BULK to ACCSAT, because ACCSAT spills more registers to global memory, facilitating the reuse of computation. The next three kernels execute 8.3% fewer instructions and achieve around a 1.11x speedup compared to CSE+BULK, since our optimized code clarifies dependencies and reduces both computation and stores. We expect a total speedup of 1.22x when equality saturation is disabled when registers spill.
@@
Figure 4 depicts the speedups of SPEC benchmarks for both OpenACC and OpenMP versions. Benchmarks with names starting with ”p-” indicate the OpenMP versions. On average, with OpenACC, CSE, CSE+SAT, CSE+BULK, and ACCSAT attain 1.06x, 1.04x, 1.08x, and 1.08x speedups, respectively on NVHPC, while with OpenMP, these attain 1.07x, 1.06x, 1.37x, and 1.47x speedups, respectively. OpenACC’s ostencil sees a 16.7% reduction in SM occupancy due to equality saturation, leading to decreased performance. CSE reduces memory loads by around 50% for olbm and polbm, resulting in improved performance. Bulk load significantly boosts the performance of memory/intensive benchmarks such as csp, bt, pcsp, and pbt, just as in NPB. ACCSAT attains 3.62x speedup in one part of pbt, which executes only one thread-block over nested loops, by eliminating 77.2% memory loads and 50.9% stores. Our optimization decreases operations and reorders memory accesses, being effective for both parallel and sequential iterations.
@@
IX Related Work
Maximizing architectural utilization is a significant challenge in HPC, requiring application code optimization. Compilers play a crucial role for enhancing performance by applying various code-generation techniques, enabling both generic and architecture-specific optimizations. Numerous programming models and state-of-the-art techniques, especially for GPUs, have been introduced to adapt codes to parallel architectures [37, 3, 38, 39, 40]. Directive-based programming models [5, 6] extend sequential languages, allowing complex scientific applications to offload loop iterations to accelerators while maintaining their structures. However, such compilers often rely on the basis of sequential code generation, thus limiting optimization opportunities to general computation [8, 7].
Several projects have explored domain-specific or architecture/specific approaches for directive-based code. The CLAW DSL [41] provides directives for grid-based algorithms, enabling target-specific optimizations such as spatial blocking while supporting OpenACC/OpenMP code generation. JACC [8] is an OpenACC runtime framework offering just-in-time kernel compilation with dynamic constant expansion. OptACC [24] performs runtime parameter searches to optimize OpenACC parallelism. CCAMP [16] interchanges OpenACC and OpenMP, optimizing parallelization for each combination of models and architecture. Barua et al. [7] develop an automated OpenACC-kernel optimizer for maximizing ILP through loop unfolding. SAFARA [42] fully utilizes register resources to facilitate array reference reuse in OpenACC kernels.
@@ IX Related Work
Since the introduction of the equality-saturation library egg [21], numerous studies have leveraged it, particularly in the context of GPU computing, for accelerating deep learning applications [43, 44, 45, 46, 47]. These works employ rewriting rules for arithmetic expressions, abstract operations, or tensor graphs to optimize convolutions, sparse tensor algebra, or whole tensor operations. Diospyros [48] synthesizes efficient DSP operations from C code using equality saturation, while Gowda et al. [49] implement a symbolic algebra system with egg for automatic parallelism assignment.
Although the initial work of equality saturation was demonstrated as a Java bytecode optimizer [9], recent works using egg focus on program synthesis and code optimization without control statements. Our work is the first to bridge the gap between user code and equality saturation by automatically extracting computation and constructing SSA information for data dependencies. This approach enables novel equality/saturation optimizations for directive-based programming, as detailed in this paper, without requiring further programmer intervention.
Several works propose innovative code optimization approaches. Ben-Nun et al. [50, 51] develop data-centric flow graphs to focus on data-oriented optimizations, such as dead memory elimination. Their representation supports collective operations and macroscopic parallelization, enabling optimizations irrespective of calculation. The MLIR framework [52] utilizes multi-level IRs for cooperative domain-specific optimizations. Ginsbach et al. [53] define code patterns for performance opportunities, replacing matched code with library or DSL implementations. Additionally, machine learning is gaining popularity for automatic compiler-optimization tuning [54, 55, 56]. Our tool employs e-graph operations to identify optimal solutions without source code analysis or abstract syntax tree transformations.
Numerous optimization techniques have been developed for GPU computing [57]. Rawat et al. [20] reorder stencil computation using heuristic algorithms to alleviate register pressure and increase SM occupancy. Software systolic arrays [58] assign global data to GPU threads, computing results by propagating and accumulating partial results through warp-level shuffles across overlapped blocks, improving data locality without shared memory usage. Hong et al. [59] utilize lightweight kernel emulation to provide auto-tuners with performance bottleneck feedback. In contrast, ACC Saturator adopts a straightforward rewriting mechanism, enabling its application to complex programs. Moreover, our method is complementary to other code optimizations.
X Conclusion